OpenStudy (anonymous):

How many moles of potassium are represented by 5.60 ✕ 10e22 potassium atoms?

OpenStudy (jhannybean):

You should know that \(1 ~\text{mole} = ~6.022 \times 10^{23}~\sf \text{particles}\)\[\sf 5.60 \times 10^{22}~K^+ \text{atoms}~ \cdot \frac{1 ~~mole ~~K^+}{6.022\times 10^{23} ~K^{+}\text{ atoms}}\]

OpenStudy (jhannybean):

Remember Avogadro's number?